OpenStudy (anonymous):

Solve: x2 – x = 2 a) x = 2, x = -1 b) x = -2, x = 1 c) x = -1, x = -2 d) x = 1, x = 2

sam (.sam.):

you need to set the left-hand side of the equation equal to 0, x^(2)-x-2=0 1*-2=-2 and 1-2=-1, so insert 1 as the right hand term of one factor and -2 as the right-hand term of the other factor. (x+1)(x-2)=0 x=-1,2